OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = −1 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (aroub):

2x + y = 1 4x + 2y = −1 y=-2x+1 4x+2(-2x+1)=-1 4x-4x+2=-1 0=-3 There is no x value as there is no solution.